# Cosmogenic Origin of KM3-230213A: Delayed Gamma-Ray Emission from A Cosmic-Ray Transient

### v1

@@ Abstract
The highest-energy cosmic neutrino detected by the ARCA detector of KM3NeT has reignited the quest to pinpoint the sources of ultrahigh-energy cosmic rays (UHECRs; $E\gtrsim 0.1$ EeV). By uncovering the associated multimessenger signals, we investigate the origin of the 220 PeV $\nu_{\mu}$ event KM3-230213A from a transient source that accelerated cosmic rays to $\sim 10$ EeV. UHECR protons that escape the source interact with the cosmic background radiation, producing a PeV-EeV cosmogenic neutrino spectrum. The secondary $e^{\pm}$ and $\gamma$-rays initiate an electromagnetic cascade, resulting in a cosmogenic $\gamma$-ray spectrum that peaks at a delayed time due to deflection of charged particles in the extragalactic magnetic field (EGMF). Our results shed light on the nature of the UHECR source for the $\nu_{\mu}$ event and provide crucial insights into the detection of multi-TeV $\gamma$-rays of cosmogenic origin from similar past cosmological transients. Using the $\gamma$-ray sensitivity of currently operating and next-generation imaging atmospheric Cherenkov telescopes, the flux and time-delay distribution can constrain the source distance. We further show that the detection of such a $\gamma$-ray signal above the background depends on the EGMF strength. Together with the non-detection of coincident spatial or temporal photon counterparts at the current epoch, this detection is the first compelling candidate for a sub-EeV cosmogenic neutrino.
@@ 1 Introduction
In this letter, we analyse the cosmogenic fluxes from a transient point source. The all-sky diffuse cosmogenic neutrino flux arises from integrating over all point sources at all redshifts. However, we assume that the transient source considered in our analysis belongs to a subdominant unresolved UHECR source population that injects protons, consistent with Fermi-LAT IGRB constraints on a proton-dominated UHECR source model. We present the prospects for detecting the associated cosmogenic $\gamma$-ray flux, peaking at a much later time, if the neutrino event originated from a UHECR interaction. We consider various source distances and examine the time-delay distributions of neutrinos and $\gamma$-rays. Our analysis constrains the source distance by the absence of coincident electromagnetic counterparts at the current epoch. This analysis also provides insight into orphan very-high-energy (VHE; $E_{\gamma}\gtrsim 0.1$ TeV) $\gamma$-ray detection as a delayed signal from a past transient and hints at the nature of such sources, which produce a neutrino flux comparable to that of the KM3-230213A event at a much earlier time.
In Sec. 2, we summarize our model considerations and the simulation setup used to analyze cosmogenic neutrino and $\gamma$-ray fluxes. We present our results in Sec. 3 and discuss the implications in Sec. 4. We draw our conclusions in Sec. 5.
2 Cosmogenic fluxes and Simulation Setup
@@
We employed the Monte Carlo simulation framework CRPropa3 (AlvesBatista:2016vpy; AlvesBatista:2022vem) to simulate the 3-D propagation and interactions of UHECRs and to track the development of the resulting electromagnetic (EM) cascades initiated by secondary $e^{\pm}$ and $\gamma$-rays. We set up a random turbulent extragalactic magnetic field (EGMF) with a Kolmogorov power spectrum over length scales of 100 kpc to 5 Mpc, and a turbulence coherence length of $\approx 1$ Mpc. We choose an optimistic value of the RMS field strength $B_{\rm RMS}\approx 10^{-5}$ nG. Fermi observations of TeV blazars constrain the EGMF to be as low as $\sim 10^{-7}$ nG (Neronov_2010). The magnetic field is periodically repeated in space to cover an arbitrary simulation volume. Due to computational cost, we calculate our cascade photon flux down to 10 GeV. Lower-energy photons are expected to arrive at higher deflection angles and are difficult to associate with the transient that produces such a neutrino event.
@@
UHECRs and secondary $e^{\pm}$ are deflected in cosmic magnetic fields, causing a delay in the cosmogenic $\gamma$-ray signal. For neutrinos, UHECR interactions with the EBL are the dominant contribution in our chosen energy range. The corresponding cosmogenic neutrino flux is normalised to the flux predicted by the KM3NeT-only fit. The per-flavor diffuse flux is $E_{\nu}^{2}\Phi_{\nu+\overline{\nu}}^{\rm 1f}=5.8\times 10^{-8}$ GeV cm-2 s-1 sr-1 assuming a $E_{\nu}^{-2}$ spectrum in the central 90% energy range of 72 PeV - 2.6 EeV (KM3NeT:2025ccp). Since the flux corresponds to a single observed event, we consider the KM3NeT-only fit for our analysis of a transient point source and convert it to point-like flux by multiplying with $4\pi$. The initial emission direction of UHECRs is collimated along our line of sight. Secondary neutrinos and $\gamma$-rays are collected on the surface of an observer sphere of radius such that a maximum deflection of $1^{\circ}$ is recorded, i.e., the observer sphere increases in size as the distance to the source increases (Elyiv_2009). The KM3NeT-ARCA event has a 99% C.L. angular uncertainty of $3^{\circ}$. However, the Fermi-LAT resolution is $0.15^{\circ}$ for photons of energy >30>\!30 GeV.
3 Results
Fig. 1 shows a schematic representation of the scenario considered here. The straight line path from the source to the observer corresponds to the case of zero time delay between neutrino and $\gamma$-ray signals. UHE protons can be deflected by the EGMF, leading to arrival delays in observed signals (Kotera:2008ae; Dermer:2008cy; Murase_2012). Neutrinos, although propagating rectilinearly once produced, may still inherit delays due to proton propagation time before neutrino production and the transverse offset between neutrino production site and the source-observer line of sight. Cosmogenic photons experience further delays due to the EM cascade, where pair production and inverse-Compton scattering lengthen the effective propagation path through successive cascade generations.
[…]
@@ 3 Results
The $\gamma$-ray flux peaks at a much later time, as shown in the subsequent Figs. 3-6. For $z=0.01$, the comoving source distance is $d_{c}\approx 133$ Mpc and the electromagnetic cascade is not fully developed. As expected for a secondary cascade, the event fraction of lower-energy (10–100 GeV) photons (blue lines) is higher at later times, while at early times the delay distribution (left panels) is comparable to that of higher-energy (100 GeV–100 TeV) photons. This general trend becomes more prominent at higher redshifts, owing to the sufficient cascade of UHE $e^{\pm}$ and $\gamma$-rays. The right panels show the $\gamma$-ray flux over the entire energy range in our analysis and for various time windows. The fluxes are obtained by normalizing the neutrino spectrum from our simulations to the KM3NeT-only fit flux level at 220 PeV, after accounting for the additional factor due to per-flavor flux. Since a significant fraction of the neutrino events arrive near $\Delta t=0$, we use the single-event KM3NeT flux for the normalization, with negligible impact on our results. We overlay the sensitivity of currently operating and upcoming telescopes, viz., Fermi-Large Area Telescope (LAT), the northern and southern arrays of Cherenkov Telescope Array Observatory (CTAO) for 50h of observation (CTAConsortium:2017dvg), Large High-Altitude Air-Shower Observatory (LHAASO) (LHAASO:2019qtb), and Southern Wide-field Gamma-ray Observatory (SWGO) (Albert:2019afb). We find that for source distances $z\lesssim 0.2$, the $\gamma$-ray spectrum is comparable to the sensitivities of the upcoming detectors, peaking at $\sim 10$ TeV and time as early as $\Delta t\approx 0.1$ Myr. Whereas, for higher source distances, the cascade is fully developed and high-energy $\gamma$-rays are further absorbed in the EBL, leading to a spectrum that peaks at $\mathcal{O}\sim 100$ GeV and at later times. Due to computational limitations, we do not show the $\gamma$-ray fluxes for shorter time intervals.
@@ 4 Discussions
Our numerical analysis incorporating 3-D deflections of UHE protons in the EGMF reveals a time delay of cosmogenic $\gamma$-rays that peaks at around $\sim 100$ kyr with respect to the light crossing time. This is consistent with the analytical estimates of the time delay due to inverse-Compton cooling and bending of the electron trajectory in the EGMF. Hence, if the origin of KM3-230213A is truly cosmogenic, the time evolution of the $\gamma$-ray spectrum shows that a detection at the current epoch is challenging. However, for similar cosmological transients in the past, the gamma-ray spectrum peaks at $\sim 10$ TeV for a source at $z\lesssim 0.2$. For a source at $\sim 0.1$, as seen in Fig. 4, the event fraction of 100 GeV – 100 TeV photons is significant at $\Delta t\sim 10^{4}$, when the neutrino event rate declines. At higher source distances, the neutrino spectrum evolves over Myr time-delay scales within the 90% central energy range of the KM3NeT event. However, at higher redshifts, the peak energy of the $\gamma$-rays shifts to lower energy, as expected for a fully developed EM cascade, and is below the detection threshold for a prudent choice of EGMF. This implies multi-TeV $\gamma$-ray and sub-EeV neutrino flux may be correlated for comoving source distances $\lesssim 500$ Mpc.
For known transient sources such as radio-loud AGNs, a significant electromagnetic counterpart is expected. Earlier studies of blazars invoking cosmogenic $\gamma$-rays indicate appreciable interactions within the sources (see, e.g., Keivani:2018rnh; MAGIC:2022gyf; Das:2022nyp) leading to a considerable hadronic component in X-rays and $\gamma$-rays. Similarly, in the case of gamma-ray bursts (GRBs), a prompt $\gamma$-ray/X-ray emission or extended multiwavelength afterglow emission is expected. The non-detection of high-energy neutrinos from GRBs by IceCube (IceCube:2014jkq; IceCube:2017amx) has severely constrained cosmic-ray acceleration during the prompt emission phase. Sub-threshold $\gamma$-ray sources, including transients, that are not included in Fermi catalogues have been proposed (Sherman:2025gir). For radio-quiet AGNs or Seyfert galaxies, reprocessed GeV-TeV $\gamma$-rays contribute to a significant X-ray emission (Kheirandish:2021wkm; Murase:2022dog). Hence, the time delay between a prompt neutrino produced near the source and its electromagnetic counterpart may pose difficulties with known transient models (Das:2025vqd). However, the proximity of the KM3-230213A event to the Galactic plane may have hindered source identification due to Galactic foreground emission.
5 Conclusions
@@
6 Acknowledgement
S.B. and N.G. thank R. Alves Batista for helpful discussions on estimating trajectory lengths with CRPropa3.

### v2

@@ Abstract
The highest-energy cosmic neutrino detected by the ARCA detector of KM3NeT has reignited the quest to pinpoint the sources of ultrahigh-energy cosmic rays (UHECRs; $E\gtrsim 0.1$ EeV). By uncovering the associated multimessenger signals, we investigate the origin of the 220 PeV $\nu_{\mu}$ event KM3-230213A from an unknown transient that accelerated cosmic rays to $\sim 10$ EeV. Unlike an astrophysical origin, where the $\nu_{\mu}$ is produced inside the source, here we consider UHECR protons that escape the source interact with the cosmic background radiation, producing a PeV-EeV cosmogenic neutrino spectrum. The secondary $e^{\pm}$ and $\gamma$-rays initiate an electromagnetic cascade, resulting in a cosmogenic $\gamma$-ray spectrum. The latter peaks at a delayed time of $\gtrsim 10^{4}$ years compared to the light travel time from the transient to observer, due to deflection of charged particles in the extragalactic magnetic field (EGMF). Our results shed light on the nature of the UHECR source for the $\nu_{\mu}$ event and provide crucial insights into the detection of multi-TeV $\gamma$-rays of cosmogenic origin from similar past cosmological transients. Using the $\gamma$-ray sensitivity of currently operating and next-generation imaging atmospheric Cherenkov telescopes, the flux and time-delay distribution can constrain the source distance. We further show that the detection of such a $\gamma$-ray signal above the background depends on the EGMF strength. Together with the non-detection of coincident spatial or temporal photon counterparts at the current epoch, this detection is the first compelling candidate for a sub-EeV cosmogenic neutrino.
@@ I Introduction
In this letter, we analyse the cosmogenic fluxes from a transient point source. Here, the unknown transient phenomenon may be a short-lived astrophysical event characterized by an episode of increased activity (and thus increased cosmic-ray luminosity) relative to the emission during the baseline quiescent state. The all-sky diffuse cosmogenic neutrino flux arises from integrating over all point sources at all redshifts. However, we assume that the transient source considered in our analysis belongs to a subdominant unresolved UHECR source population that injects protons, consistent with Fermi-LAT IGRB constraints on a proton-dominated UHECR source model. We present the prospects for detecting the associated cosmogenic $\gamma$-ray flux, peaking at a much later time, if the neutrino event originated from a UHECR interaction. We consider various source distances and examine the time-delay distributions of neutrinos and $\gamma$-rays. Our analysis constrains the source distance by the absence of coincident electromagnetic counterparts at the current epoch. This analysis also provides insight into orphan very-high-energy (VHE; $E_{\gamma}\gtrsim 0.1$ TeV) $\gamma$-ray detection as a delayed signal from a past transient and hints at the nature of such sources, which produce a neutrino flux comparable to that of the KM3-230213A event at a much earlier time.
In Sec. II, we summarize our model considerations and the simulation setup used to analyze cosmogenic neutrino and $\gamma$-ray fluxes. We present our results in Sec. III and discuss the implications in Sec. IV. We draw our conclusions in Sec. V.
II Cosmogenic fluxes and Simulation Setup
@@
We employed the Monte Carlo simulation framework CRPropa3 (AlvesBatista:2016vpy; AlvesBatista:2022vem) to simulate the 3-D propagation and interactions of UHECRs and to track the development of the resulting electromagnetic (EM) cascades initiated by secondary $e^{\pm}$ and $\gamma$-rays. We set up a random turbulent extragalactic magnetic field (EGMF) with a Kolmogorov power spectrum over length scales of 100 kpc to 5 Mpc, and a turbulence coherence length of $\approx 1$ Mpc. We choose an optimistic value of the RMS field strength $B_{\rm RMS}\approx 10^{-5}$ nG. Fermi observations of TeV blazars constrain the EGMF to be as low as $\sim 10^{-7}$ nG (Neronov_2010). The magnetic field is periodically repeated in space to cover an arbitrary simulation volume. Due to computational cost, we calculate our cascade photon flux down to 10 GeV. Lower-energy photons are expected to arrive at higher deflection angles and are difficult to associate with the transient that produces such a neutrino event. We calculate the time delay of the cosmogenic $\gamma$ ray fluxes with respect to the light travel time from the source to the observer. This is much longer than that due to $\gamma$-ray induced cascade resulting from an astrophysical origin, as shown in Fang:2025nzg.
@@
UHECRs and secondary $e^{\pm}$ are deflected in cosmic magnetic fields, causing a delay in the cosmogenic $\gamma$-ray signal. For neutrinos, UHECR interactions with the EBL are the dominant contribution in our chosen energy range. The corresponding cosmogenic neutrino flux is normalized to the flux predicted by the KM3NeT-only fit. The per-flavor diffuse flux is $E_{\nu}^{2}\Phi_{\nu+\overline{\nu}}^{\rm 1f}=5.8\times 10^{-8}$ GeV cm-2 s-1 sr-1 assuming a $E_{\nu}^{-2}$ spectrum in the central 90% energy range of 72 PeV - 2.6 EeV (KM3NeT:2025ccp). Since the flux corresponds to a single observed event, we consider the KM3NeT-only fit for our analysis of a transient point source and convert it to point-like flux by multiplying by $4\pi$. The initial emission direction of UHECRs is collimated along our line of sight. Secondary neutrinos and $\gamma$-rays are collected on the surface of an observer sphere of radius such that a maximum deflection of $1^{\circ}$ is recorded, i.e., the observer sphere increases in size as the distance to the source increases (Elyiv_2009). The KM3NeT-ARCA event has a 99% C.L. angular uncertainty of $3^{\circ}$. However, the Fermi-LAT resolution is $0.15^{\circ}$ for photons of energy >30>\!30 GeV.
III Results
Fig. 1 shows a schematic representation of the scenario considered here. The straight line path from the source to the observer corresponds to the case of zero time delay between neutrino and $\gamma$-ray signals. UHE protons can be deflected by the EGMF, leading to arrival delays in observed signals (Kotera:2008ae; Dermer:2008cy; Murase_2012). Neutrinos, although propagating rectilinearly once produced, may still inherit delays due to proton propagation time before neutrino production and the transverse offset between neutrino production site and the source-observer line of sight. Cosmogenic photons experience additional delays due to the EM cascade, in which pair production and inverse-Compton scattering lengthen the effective path length through successive cascade generations.
The additional delay in the cosmogenic photon spectrum depends on the deflection of secondary electrons during the inverse-Compton energy loss time of $e^{\pm}$ pair on CMB, $t_{\rm IC}=3m_{e}c/4\gamma_{e}\sigma_{T}u_{\rm CMB}\approx 0.5\times 10^{13}(\gamma_{e}/10^{7})^{-1}$ s at $z=0.1$ (Razzaque:2004cx). The deflection angle of the pairs can be approximated as $\theta_{B}\approx ct_{\rm IC}/R_{L}\approx 8.8\times 10^{-2}(\gamma_{e}/10^{7})^{-2}(B_{\rm rms}/10^{-14}\text{\ G})$, where $R_{L}=\gamma_{e}m_{e}c^{2}/eB_{\rm rms}$ is the Larmor radius of the pairs in the EGMF. For a pencil beam, this corresponds to $\Delta t_{B}=t_{\rm IC}\theta_{B}^{2}/2\approx 600\text{\ yrs}\times(\gamma_{e}/10^{7})^{-5}(B_{\rm rms}/10^{-14}\text{\ G})^{2}$. However, in reality, the $\gamma\gamma$ pair production length plays a crucial role, and $e^{\pm}$ pairs have a finite spreading. Hence, the time delay could be written as $\Delta t=\lambda_{\rm\gamma\gamma}\theta_{B}^{2}/2c\approx 3.8\times 10^{5}$ yrs $\times(\gamma_{e}/10^{7})^{-4}\times(n_{\rm CIB}/0.1\text{\ cm}^{-3})^{-1}\times(B_{\rm rms}/10^{-14}\text{\ G})^{2}$ at $z\approx 0.1$, where $n_{\rm CIB}$ and $\lambda_{\rm\gamma\gamma}$ are the infrared photon number density and the energy loss length due to pair-production (Murase:2008pe). Whereas neutrinos, once produced, travel along a rectilinear path undeflected by the magnetic field and unattenuated by cosmic background radiation.
@@ III Results
The $\gamma$-ray flux peaks at a much later time, as shown in the subsequent Figs. 3-6. For $z=0.01$, the comoving source distance is $d_{c}\approx 133$ Mpc and the electromagnetic cascade is not fully developed. As expected for a secondary cascade, the event fraction of lower-energy (10–100 GeV) photons (blue lines) is higher at later times, while at early times the delay distribution (left panels) is comparable to that of higher-energy (100 GeV–100 TeV) photons. This general trend becomes more prominent at higher redshifts, owing to the sufficient cascade of UHE $e^{\pm}$ and $\gamma$-rays. The right panels show the $\gamma$-ray flux over the entire energy range in our analysis and for various time windows. The fluxes are obtained by normalizing the neutrino spectrum from our simulations to the KM3NeT-only fit flux level at 220 PeV, after accounting for the additional factor due to per-flavor flux after neutrino oscillation. Since a significant fraction of the neutrino events arrive near $\Delta t=0$, we use the single-event KM3NeT flux for the normalization, with negligible impact on our results. We overlay the sensitivity of currently operating and upcoming telescopes, viz., Fermi-Large Area Telescope (LAT), the northern and southern arrays of Cherenkov Telescope Array Observatory (CTAO) for 50h of observation (CTAConsortium:2017dvg), Large High-Altitude Air-Shower Observatory (LHAASO) (LHAASO:2019qtb), and Southern Wide-field Gamma-ray Observatory (SWGO) (Albert:2019afb). We find that for source distances $z\lesssim 0.2$, the $\gamma$-ray spectrum is comparable to the sensitivities of the upcoming detectors, peaking at $\sim 10$ TeV and time as early as $\Delta t\approx 0.1$ Myr. Whereas, for higher source distances, the cascade is fully developed and high-energy $\gamma$-rays are further absorbed in the EBL, leading to a spectrum that peaks at $\mathcal{O}\sim 100$ GeV and at later times. Due to computational limitations, we do not show the $\gamma$-ray fluxes for shorter time intervals.
@@
The sensitivities of current and next-generation facilities are essential for probing such past transients. For cosmological events of this type that occurred $\sim 10^{4}\text{--}10^{6}\,\mathrm{yr}$ ago, the delayed gamma-ray emission may be detectable at the current epoch. While the accompanying neutrino burst would have passed, the $\gamma$-ray component can be distinguished from the diffuse background due to its hard spectrum at very high energies ($\mathcal{O}\sim 10\,\mathrm{TeV}$). Moreover, the peak energy, set by the extragalactic magnetic field, provides additional diagnostic power for source identification. The non-detection of such fluxes by CTAO would imply a transient distance $\gtrsim 500\,\mathrm{Mpc}$. These predictions open a new window to study UHECR production by past cosmological transients.
@@ IV Discussions
Our numerical analysis incorporating 3-D deflections of UHE protons in the EGMF reveals a time delay of cosmogenic $\gamma$-rays that peaks at around $\sim 100$ kyr with respect to the light crossing time. This is consistent with the analytical estimates of the time delay due to inverse-Compton cooling and bending of the electron trajectory in the EGMF. Hence, if the origin of KM3-230213A is truly cosmogenic, the time evolution of the $\gamma$-ray spectrum shows that a detection at the current epoch is challenging. However, for similar cosmological transients in the past, the gamma-ray spectrum peaks at $\sim 10$ TeV for a source at $z\lesssim 0.2$. For a source at $\sim 0.1$, as seen in Fig. 4, the event fraction of 100 GeV – 100 TeV photons is significant at $\Delta t\sim 10^{4}$, when the neutrino event rate declines. At higher source distances, the neutrino spectrum evolves over Myr time-delay scales within the 90% central energy range of the KM3NeT event. However, at higher redshifts, the peak energy of the $\gamma$-rays shifts to lower energy, as expected for a fully developed EM cascade, and is below the detection threshold for a prudent choice of EGMF. This implies multi-TeV $\gamma$-ray and sub-EeV neutrino flux may be correlated for comoving source distances $\lesssim 500$ Mpc. The effects of the Galactic magnetic field can be neglected here, since nearly all secondaries are produced at distances from Earth much larger than the virial radius of our Galaxy and well beyond the observer size in our simulations.
For known transient sources such as radio-loud AGNs, a significant electromagnetic counterpart is expected. Earlier studies of blazars invoking cosmogenic $\gamma$-rays indicate appreciable interactions within the sources (see, e.g., Keivani:2018rnh; MAGIC:2022gyf; Das:2022nyp) leading to a considerable hadronic component in X-rays and $\gamma$-rays. Similarly, in the case of gamma-ray bursts (GRBs), a prompt $\gamma$-ray/X-ray emission or extended multiwavelength afterglow emission is expected. The non-detection of high-energy neutrinos from GRBs by IceCube (IceCube:2014jkq; IceCube:2017amx) has severely constrained cosmic-ray acceleration during the prompt emission phase. Sub-threshold $\gamma$-ray sources, including transients, that are not included in Fermi catalogues have been proposed (Sherman:2025gir). For radio-quiet AGNs or Seyfert galaxies, reprocessed GeV-TeV $\gamma$-rays contribute to a significant X-ray emission (Kheirandish:2021wkm; Murase:2022dog). Hence, the time delay between a prompt neutrino produced near the source and its electromagnetic counterpart may pose difficulties with known transient models (Das:2025vqd). Thus, the exact nature of the transient is difficult to derive without an electromagnetic counterpart. However, the proximity of the KM3-230213A event to the Galactic plane may have hindered source identification due to Galactic foreground emission.
V Conclusions
@@
VI Acknowledgement
S.B. and N.G. thank R. Alves Batista for helpful discussions on estimating trajectory lengths with CRPropa3. We thank the anonymous referee for valuable comments to improve the manuscript.